OpenStudy (dls):

A projectile is fired with kinetic energy 1kJ. If its range is maximum,what will be its kinetic energy at the highest point?

OpenStudy (dls):

@experimentX @hartnn

OpenStudy (anonymous):

Find potential energy first,

OpenStudy (anonymous):

For maximum range, if there's no air friction, the angle must be 45 degrees.

OpenStudy (dls):

Rmax=u^2/g

OpenStudy (anonymous):

what's u?

OpenStudy (dls):

initial velocity

OpenStudy (anonymous):

Try this: h max=v0y^2/2g h max=(v0*sin45)^2/2g What's the potential energy at h max?

OpenStudy (anonymous):

i hope the answer is 1/2 kj

OpenStudy (dls):

yeah..500J..thanks!

OpenStudy (anonymous):

yo i got it u want explanation ..................